In mid-2009, Rite Aid had CCC-rated, 6-year bonds outstanding with a yield to maturity of 17.3%. At the time, similar maturitplease show all work
Answer #1

a) The formula used to determine the Probability of default:

Probability of default as per market risk premium:

P= S * (1+r)/(1-RR) , where:
S = spread = 0.05*0.31 = 1.55%
R = risk-free rate = 3%
RR = expected recovery rate ad default = 40% (since expected loss rate is 40%)

P = 0.0155 * (1.03)/(0.6) = 2.660833%

Probability of default as per Yield to maturity of bond:

P= S * (1+r)/(1-RR) , where:
S = spread = 17.3% - 3% = 14.3%
R = risk-free rate = 3%
RR = expected recovery rate ad default = 40% (since expected loss rate is 40%)

P = 0.143 * (1.03)/(0.6) = 24.54833%

b) The formula used to determine the Probability of default:

P= S * (1+r)/(1-RR) , where:
S = spread = 7.1% - 1.5% = 5.6%
R = risk-free rate = 3%
RR = expected recovery rate ad default = 40% (since expected loss rate is 40%)

P = 0.056 * (1.03)/(0.6) = 9.6133%

c) Beta = Covariance of Market returns with Stock Returns / Variance of Market returns. The following table computes Beta:

2010 2011 2012 2013 2014
Market Returns 10% -5% 10% 7% 3%
Stock Returns 15% -9% 20% 12% 2%
Premium 5% -4% 10% 5% -1%
Covariance 0.0057
Variance 0.00316
Beta 1.80379747
Average market returns 5%

Expected Equity return = risk-free rate + market portfolio premium * Equity Beta

Here,
the risk-free rate has to be assumed. Let us say it is 3%

Market portfolio premium = Average market return - risk-free rate = 5%-3% = 2%

Expected equity return = 3% + 2% * 1.80379747 = 6.6075949%

d) WACC = (VE/V ​× Re)+((VD/V))×Rd×(1−Tc))
where,
VE=Market value of the firm’s equity= $300m
VD=Market value of the firm’s debt = $100m
V=VE+VD = $400m
Re=Cost of equity = 6.6075949%, as given in part c of question
Rd=Cost of debt = 7.1%, as given in Part b
Tc=Corporate tax rate​ = 25%

WACC = (0.75 ​× 6.6075949%)+((0.25))×7.1%×(0.75)) = 6.2869462%